### v1

# Selective labeling with false discovery rate control

Abstract
Obtaining high-quality labels for large datasets is expensive, requiring massive annotations from human experts. While AI models offer a cost-effective alternative by predicting labels, their label quality is compromised by the unavoidable labeling errors. Existing methods mitigate this issue through selective labeling, where AI labels a subset and human labels the remainder. However, these methods lack theoretical guarantees on the quality of AI-assigned labels, often resulting in unacceptably high labeling error within the AI-labeled subset. To address this, we introduce Conformal Labeling, a novel method to identify instances where AI predictions can be provably trusted. This is achieved by controlling the false discovery rate (FDR), the proportion of incorrect labels within the selected subset. In particular, we construct a conformal $p$-value for each test instance by comparing AI models’ predicted confidence to those of calibration instances mislabeled by AI models. Then, we select test instances whose $p$-values are below a data-dependent threshold, certifying AI models’ predictions as trustworthy. We provide theoretical guarantees that Conformal Labeling controls the FDR below the nominal level, ensuring that a predefined fraction of AI-assigned labels is correct on average. Extensive experiments demonstrate that our method achieves tight FDR control with high power across various tasks, including image and text labeling, and LLM QA.
1 Introduction
Large-scale, high-quality labeled data is crucial for the machine learning pipelines (Deng et al., 2009). While experts could provide high-quality labels for moderately sized datasets, the growing size of modern datasets has made this approach prohibitively expensive. AI models offer a cost-effective alternative by predicting labels, bypassing the need for human experts. However, AI models are prone to labeling error (Northcutt et al., 2021; Tan et al., 2024). For example, empirical evidence demonstrates that even state-of-the-art LLMs exhibit high labeling error when used for text annotation Baumann et al. (2025). The labeling error inherent to AI models significantly compromises their label quality, hindering the deployment of AI labeling for production. To balance the trade-off between labeling cost and error, selective labeling has been a promising solution (Geifman & El-Yaniv, 2017; Wang et al., 2023) by combining AI predictions with expert annotations.
Prior work on selective labeling primarily designed heuristic methods (Wang et al., 2021; Bernhardt et al., 2022; Wang et al., 2024a) that rely on AI models to label high-confidence instances while deferring the rest to human experts. However, these methods do not provide any theoretical guarantees on label quality. To address this, a recent work (Candès et al., 2025) proposes probably approximately correct (PAC) labeling, which guarantees that the overall labeling error is controlled with high probability. They achieve the guarantee by counterbalancing the error of AI labels with the zero error of expert labels. While the guarantee is theoretically appealing, the labeling error of AI models can be unacceptably high, even reaching 100%, while the overall labeling error is controlled. These limitations motivate us to investigate how to provably guarantee the quality of AI-assigned labels.
In this work, we propose Conformal Labeling, which formulates the labeling problem as multiple hypotheses testing. This method allows us to select a subset of the unlabeled dataset for AI labeling and ensure the quality of AI-assigned labels by providing a rigorous guarantee on the false discovery rate (FDR). In particular, given a labeled calibration dataset, we compute conformal $p$-values for test instances through a rank-based comparison of their predicted confidence against those of calibration instances where AI predictions are incorrect. This ensures that $p$-values for mislabeled instances stochastically dominate the uniform distribution on [0, 1], while $p$-values for correctly labeled instances are concentrated near zero, allowing selection procedures to distinguish between them easily. Then, we select test instances whose $p$-values are less than or equal to a data-driven threshold, which is delicately set with the calibration dataset to control FDR at the desired level. We further theoretically prove that Conformal Labeling achieves valid FDR control under mild assumptions, ensuring that the expected proportion of incorrect labels in the selected subset is below the desired level. While the label quality of existing methods depends heavily on model performance, Conformal Labeling guarantees label quality regardless of the underlying model’s performance by strictly controlling the FDR.
To validate our method, we conduct extensive experiments on various labeling tasks, including image labeling (ImageNet (Deng et al., 2009), ImageNet-V2 (Recht et al., 2019)), text labeling (stance on global warming (Luo et al., 2021), and misinformation (Gabriel et al., 2022)), and LLM QA (MedMCQA (Pal et al., 2022), MMLU (Hendrycks et al., 2021), MMLU-Pro (Wang et al., 2024a)) tasks. The results demonstrate that Conformal Labeling achieves high power with controlled FDR, indicating that AI models can label a large proportion of data with high quality. For example, Conformal Labeling can label 58.67% of the ImageNet dataset with ResNet-34 (He et al., 2016), while keeping the FDR below 10%. In comparison, a naive approach of using AI-assigned labels for the entire dataset results in a labeling error of over 25%. Moreover, through comprehensive ablation studies, we validate that Conformal Labeling is robust to the size of calibration datasets, and more powerful models enable better selection results.
We summarize our contributions as follows:
  • We propose Conformal Labeling, a novel method for identifying instances where AI predictions could be provably trusted. Regardless of AI models’ performance, Conformal Labeling guarantees the quality of AI-assigned labels by strictly controlling the FDR.
  • We theoretically prove that Conformal Labeling provides a strict quality guarantee for AI-assigned labels: it achieves valid FDR control, ensuring the expected proportion of incorrect labels is below a user-specified level.
  • We empirically show that Conformal Labeling significantly reduces the labeling cost while tightly controlling the FDR, through extensive experiments conducted on image labeling, text labeling, and LLM QA tasks with various models.
2 Preliminaries
Problem setup.
In this work, we study the problem of identifying instances where AI predictions can be provably trusted. Here, we give a formulation of multi-class classification as an example. Let $\mathcal{X}$ denote the feature space and $\mathcal{Y}=\{1,\ldots,K\}$ denote the label space. The test dataset $\mathcal{D}_{\mathrm{test}}=\{X_{j}\}_{j=1}^{m}$ consists of $m$ data instances, sampled i.i.d. from a data distribution $\mathcal{P}_{\mathcal{X}}$. We denote the unseen ground-truth labels of instance $X_{j}$ as $Y_{j}$. Besides, we consider a pre-trained AI model $f:\mathcal{X}\to\mathbb{R}^{|\mathcal{Y}|}$ used to generate labels for the test dataset. For a given test instance $X$, the AI model predicts the label with the largest estimated probability $\hat{Y}=\arg\max_{y\in\mathcal{Y}}f_{y}(X)$, where $f_{y}(X)$ denotes the estimated class probability for class $y\in\mathcal{Y}$.
Since AI models are typically prone to labeling errors, we aim to select a large subset from the test dataset $\mathcal{D}_{\mathrm{label}}$ to control the portion of incorrect labels. Formally, our goal is to identify the largest subset of indices $\mathcal{R}\subseteq\{1,\cdots,m\}$ that controls the false discovery rate (FDR), defined as below:
$\text{FDR}=\mathbb{E}\left[\frac{|\mathcal{R}\cap\mathcal{H}_{0}|}{\max(|\mathcal{R}|,1)}\right],$
where $H_{0}=\{j\in\{1,\cdots,m\}:Y_{j}\neq\hat{Y}_{j}\}$ is the set of indices with incorrect predictions. For notation shorthand, we denote $m]=\{1,\cdots,m\}$ in the following. The FDR metric measures the expected proportion of mislabeled samples within the AI-labeled subset, illustrating the quality of AI-assigned labels by explicitly bounding the fraction of incorrect labels.
In addition to FDR control, we also expect AI models to label as many test instances as possible, which corresponds to maximizing power:
$\text{Power}=\mathbb{E}\left[\frac{|\mathcal{R}\cap\mathcal{H}_{1}|}{\max(|\mathcal{H}_{1}|,1)}\right],$ (1)
[…]
In this work, we assume access to a small labeled calibration dataset $\mathcal{D}_{\mathrm{cal}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$. For convenience, we denote the test dataset as $\mathcal{D}_{\mathrm{test}}=\{(X_{j},Y_{j})\}_{j=n+1}^{n+m}$, where $Y_{j}$ is not observed. Since the labeling cost of a small dataset by human annotators is typically affordable, this assumption is practical in the real world and is also adopted in prior work (Candès et al., 2025). Besides, we assume that examples of the test and calibration datasets are both drawn i.i.d. from the joint distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$, a common setting in selective labeling (Jung et al., 2024; Candès et al., 2025).
Selective labeling methods and their limitations.
Ensuring high-quality labels while reducing annotation costs has motivated extensive research on selective labeling. Prior work on selective labeling primarily focused on heuristic methods. For example, some studies design collaborative annotation frameworks that combine expert labels with LLM labels to streamline the annotation process (Li et al., 2023; Kim et al., 2024). Others propose domain-specific methods, such as meta-learning strategies for medical image labeling (Vrabac et al., 2022), annotation frameworks for text data (Duan & Lalor, 2023), and human–AI collaborative systems for object detection (Zhang et al., 2025). Although these heuristic approaches effectively reduce annotation costs, they lack formal guarantees on the label quality, which can result in unreliable labels when AI models perform poorly.
Probably approximately correct (PAC) labeling (Candès et al., 2025) addresses this limitation by providing a theoretical guarantee: the overall labeling error across the dataset is controlled with high probability. At its core, PAC labeling strategically collects zero-error expert labels for instances where the AI model exhibits the highest uncertainty, while relying on potentially imperfect AI predictions for more certain instances. This strategic allocation ensures that the dataset’s overall labeling error is small, as the zero-error expert annotations effectively counterbalance the error introduced by the AI-assigned labels. However, because the guarantee only applies at the aggregate dataset level, the subset labeled by AI may exhibit high labeling error. This limitation underscores a critical gap: existing selective labeling methods cannot ensure the quality of AI-assigned labels, hindering their reliable deployment in real-world applications. To address this, we explore methods to provably guarantee the quality of AI-assigned labels in selective labeling.
3 Method
3.1 Conformal Labeling
Our previous section shows that existing methods cannot guarantee the quality of AI-assigned labels. To address this limitation, we propose Conformal Labeling, which identifies instances where AI predictions can be provably trusted by controlling the FDR. Our approach is composed of three primary steps: quantifying uncertainty, constructing conformal $p$-values, and thresholding.
Uncertainty quantification.
Our approach builds on a key insight: we should select instances where the model exhibits high confidence in its predictions. To quantify the model confidence, we define an uncertainty score $\mathcal{S}:\mathcal{X}\to\mathbb{R}$, where a higher value indicates greater model uncertainty. We note that in the conformal inference framework, this score function is also known as the non-conformity score function. For example, we employ $\mathcal{S}(X)=1-\max_{y\in\mathcal{Y}}f_{y}(X)$ as our uncertainty score function (Hendrycks & Gimpel, 2016). This score is a valid measure of uncertainty, since prior works establish that misclassified samples generally receive lower probability scores (i.e., $\max_{y\in\mathcal{Y}}f_{y}(X)$) than correctly classified ones (Hendrycks & Gimpel, 2016; Tu et al., 2024).
Statistical guarantee via conformal $p$-value.
To provide a statistical guarantee, we reformulate our problem as the following multiple hypothesis testing problem:
$H_{j}^{0}:Y_{n+j}\neq\hat{Y}_{n+j}\quad\text{v.s.}\quad H_{j}^{1}:Y_{n+j}=\hat{Y}_{n+j},\quad\forall j=1,\cdots,m$
Rejecting the null hypothesis $H_{j}^{0}$ indicates that $(X_{n+j},\hat{Y}_{n+j})$ should be included in the subset, as it is deemed to be classified correctly. To construct the selection subset, we employ conformal p-value, which builds upon conformal inference framework (Vovk et al., 1999; 2005). The underlying intuition is that: a test instance $X$ is likely a misclassification if its uncertainty score $\mathcal{S}(X)$ is generally larger than the scores of instances that are known to be misclassified. Leveraging this idea, conformal $p$-value is computed through a rank-based comparison of $\mathcal{S}(X)$ against uncertainty scores of misclassified instances.
Algorithm 1 Conformal Labeling
1:Mislabeled Calibration set $\mathcal{D}^{0}_{\mathrm{cal}}=\{(X_{i},Y_{i})\}_{i=1}^{n_{0}}$, test instances $\{X_{n+j}\}_{j=1}^{m}$, pre-trained classifier $f$, calibration set size $n=|\mathcal{D}_{\mathrm{cal}}|$, FDR target $\alpha\in(0,1)$.
2:# 1. Compute uncertainty scores
3:for $i=1,\dots,n+m$ do
4: Compute $\mathcal{S}_{i}:=1-\max_{y\in\mathcal{Y}}f_{y}(X_{i})$.
5:end for
6:# 2. Construct conformal p-values
7:for $j=1,\dots,m$ do
8: Construct $\hat{p}_{j}$ according to equation 2.
9:end for
10:# 3. Thresholding
11:Compute $j^{*}=\max\left\{j:\hat{p}_{(j)}\leq\frac{\alpha j(n+1)}{m(n_{0}+1)}\right\}$, where $\hat{p}_{(j)}$ is the $j$-th smallest p-value.
12:Output: $\mathcal{R}=\{j:\hat{p}_{j}\leq\hat{p}_{(j^{*})}\}$.
Formally, for the calibration dataset $\mathcal{D}_{cal}$, we identify the subset $\mathcal{D}_{cal}^{0}\subseteq\mathcal{D}_{cal}$ where instances are misclassified by the AI model. For simplicity, we denote $\mathcal{D}_{cal}^{0}=\{(X_{i},Y_{i})\}_{i=1}^{n_{0}}$, and thus $Y_{i}\neq\hat{Y}_{i}$ for $i=1,\cdots,n_{0}$. We compute the uncertainty scores for the entire dataset $\{(X_{i},Y_{i})\}_{i=1}^{n+m}$: $\mathcal{S}_{i}=1-\max_{y\in\mathcal{Y}}f_{y}(X_{i})$. Then, the conformal $p$-value for the instance $X_{n+j}$ is computed by
$\hat{p}_{j}=\frac{\sum_{i=1}^{n_{0}}\mathbf{1}\{\mathcal{S}_{i}<\mathcal{S}_{n+j}\}+\left(1+\sum_{i=1}^{n_{0}}\mathbf{1}\{\mathcal{S}_{i}=\mathcal{S}_{n+j}\}\right)\cdot U_{j}}{n_{0}+1},$ (2)
where $U_{j}\sim\mathrm{Uniform}[0,1]$ are i.i.d. uniform random variable to randomize over ties when $S_{n+j}$ equals some $S_{i}$, ensuring a continuous conformal $p$-value. Here, $\hat{p}_{j}$ quantifies how extreme the uncertainty of $X_{n+j}$ is compared to the scores of misclassified instances, with a small $\hat{p}_{j}$ providing strong statistical evidence for correct prediction.
Standard results from conformal inference establish the validity of conformal $p$-value in practice: if the instance $X_{n+j}$ is misclassified, then the conformal $p$-value stochastically dominates the uniform distribution on $[0,1]$ (Bates et al., 2023; Jin & Candès, 2023): $\mathbb{P}\{\hat{p}_{j}\leq\alpha\mid H_{j}^{0}\text{ is true}\}\leq\alpha.$ This property indicates that the conformal $p$-values for misclassified instances are biased to be high, which allows us to set a threshold to flag potential correct instances, while controlling the FDR.
Thresholding.
After obtaining conformal $p$-values for test instances, we apply a thresholding rule inspired by the Benjamini–Hochberg (BH) procedure (Benjamini & Hochberg, 1995) to select a maximal subset $\mathcal{R}$ for AI labeling while controlling the FDR at level $\alpha$. The key idea is that we gradually increase the acceptance threshold until including additional samples would risk exceeding the desired labeling error. In particular, let $p_{(1)}\leq\ldots\leq p_{(m)}$ denote the ordered statistics of the $p$-values; the rejection set of our selection procedure applied to the conformal $p$-values is $\mathcal{R}=\{j:\hat{p}_{j}\leq\hat{p}_{(j^{*})}\}$, where
$j^{*}=\max\left\{j:\hat{p}_{(j)}\leq\frac{\alpha j(n+1)}{m(n_{0}+1)}\right\},$
[…]
Theoretical results.
We now provide a theoretical guarantee for Conformal Labeling. In the following theorem, we establish that Algorithm 1 controls the FDR below the desired level $\alpha$. This theorem ensures that AI predictions can be provably trusted in the selected subset, as the expected proportion of incorrect labels is controlled below $\alpha$. The proof is provided in Appendix B.2.
Theorem 3.1.
Suppose calibration samples $(X_{i},Y_{i})_{i=1}^{n}$ and test samples $(X_{n+j},Y_{n+j})_{j=1}^{m}$ are i.i.d. Let $\alpha\in(0,1)$ be the target FDR level, and suppose the selection set $\mathcal{R}$ is determined by Algorithm 1 applied at the target FDR level $\alpha$. Define $p=\mathbb{E}[H_{j}^{0}]$, the probability that a test sample $(X_{n+j},Y_{n+j})$ is incorrectly predicted. Then, the FDR of the selection set $\mathcal{R}$ satisfies:
$\text{FDR}\leq\left[1-(1-p)^{n+1}\right]\alpha\leq\alpha.$
3.2 Choice of uncertainty score
In the above analysis, we establish that Conformal Labeling controls the FDR below the desired level. However, this guarantee alone is insufficient: a trivial procedure that simply labels nothing would achieve a perfect FDR of 0, yet offer no practical value. This highlights the need to also evaluate the method’s statistical power, which measures the method’s ability to identify as many correctly labeled instances as possible (see Eq. (1)).
As shown in prior work (Jin & Candès, 2023; Gui et al., 2024; Bai et al., 2025b), the statistical power of this method depends on the quality of the uncertainty score. In particular, a score that better separates correct from incorrect predictions directly increases statistical power (see Proposition 7 of Jin & Candès (2023)). To deliver practical recommendations, we empirically compare several uncertainty scores by visualizing their resulting $p$-value distributions and measuring the final statistical power employed with our method. We utilize a pre-trained ResNet-34 model on the ImageNet dataset, with three uncertainty scores: maximum softmax probability (MSP) (Hendrycks & Gimpel, 2016), energy score (Liu et al., 2020), and DOCTOR-$\alpha$ score ($D_{\alpha}$) (Granese et al., 2021). We give an overview of these score functions in Appendix C. The results in Figure 1 show that both MSP and $D_{\alpha}$ score provide a clear distinction between correct and incorrect predictions, thus achieving high statistical power. However, the energy score fails to provide this separation, leading to low power. Given the comparable power performance of MSP and $D_{\alpha}$ score, we will use the more computationally efficient MSP in our main experiments.
(a) MSP
(b) Energy
(c) $D_{\alpha}$
(d) Power comparison
Figure 1: Empirical distributions and power (employed with our method) of conformal $p$-values under different uncertainty scores. The experiments include maximum softmax probability (MSP), energy, and DOCTOR-$\alpha$ score ($D_{\alpha}$). The experiments are conducted on ImageNet with ResNet-34. The results show that both MSP and $D_{\alpha}$ score create a clear distinction between correct and incorrect predictions, thus achieving high statistical power. However, the energy score fails to provide this separation, leading to low power.
4 Experiments
In this section, we evaluate the effectiveness of Conformal Labeling on image labeling, text labeling, and LLM QA tasks with various models. We find that it achieves tight FDR control and high power, indicating that AI models can label a large proportion of data with high quality. We also conduct comprehensive ablation studies to provide practical guidance for applying our method.
4.1 Experimental setup
Tasks and datasets.
We evaluate the effectiveness of Conformal Labeling across three labeling tasks, including image labeling, text labeling, and LLM question answering (QA). In appendix G, we also demonstrate how to extend Conformal Labeling to regression tasks. For the LLM QA task, the goal is to identify subsets of questions that LLMs can answer correctly. We employ common benchmark datasets for evaluations in each labeling task. For image classification, we use ImageNet (Deng et al., 2009) and its variant ImageNet-V2 datasets (Recht et al., 2019). For text labeling, we adopt two benchmark datasets. The first is Stance on Global Warming (Luo et al., 2021), which provides annotations ($Y_{i}\in\{\text{agree},\text{neutral},\text{disagree}\}$) to judge whether a headline agrees that global warming is a serious concern. The second is Misinformation (Gabriel et al., 2022), which contains binary annotations ($Y_{i}\in\{\text{misinfo},\text{real}\}$) for identifying whether a given text contains misinformation. For the LLM QA task, we evaluate our method on MedMCQA (Pal et al., 2022), MMLU (Hendrycks et al., 2021), and MMLU-Pro (Wang et al., 2024b) datasets.
Table 1: Performance of Conformal Labeling on three labeling tasks. $\uparrow$ indicates that a larger value is better. We evaluate on Image Labeling (ImageNet, ImageNet-V2), text Labeling (Stance on global warming, Misinformation), and LLM QA (MedMCQA, MMLU) tasks. We report results for Conformal Labeling at $\alpha=0.1$ and compare Conformal Labeling against two baselines: (i) a naive approach of labeling instances whose uncertainty score $\mathcal{S}_{n+j}\leq 0.1$ with AI models, and (ii) labeling the entire dataset with AI models. The results show that Conformal Labeling consistently achieves tighter FDR control across all datasets and models compared with the baseline.
Task Dataset Model Conformal Labeling ($\alpha$= 0.1) Naive ($\mathcal{S}\leq 0.1$) AI only
FDR % Power %($\uparrow$) Ratio %($\uparrow$) FDR % Power %($\uparrow$) Ratio %($\uparrow$) Error %($\downarrow$)
Image ImageNet ResNet-34 9.97 80.01 58.67 4.79 63.09 43.71 26.71
DenseNet-161 9.99 85.03 65.56 5.58 72.08 52.98 22.89
ResNeXt50 10.00 86.06 66.83 6.08 75.57 56.21 22.39
CLIP-VIT-B/32 9.98 46.04 27.47 5.53 28.65 16.28 40.35
ImageNet-V2 ResNet-34 10.00 61.99 37.87 7.25 56.03 33.15 39.03
DenseNet-161 9.83 66.67 43.39 9.39 65.63 42.45 34.88
ResNeXt50 9.95 67.86 44.75 10.17 68.71 45.38 34.08
CLIP-VIT-B/32 9.96 34.56 18.10 7.86 25.83 13.18 47.78
Text Stance Llama-3.1-8B-Instruct 9.77 19.02 9.42 24.43 54.94 31.39 52.04
Qwen3-32B 9.56 10.70 7.25 14.42 16.64 11.11 36.52
Qwen2.5-72B-Instruct 9.71 26.97 17.84 26.27 74.49 59.01 35.09
Misinformation Llama-3.1-8B-Instruct 9.88 7.31 5.81 18.38 66.17 55.25 24.28
Qwen3-32B 9.91 49.21 37.48 10.77 52.27 40.03 24.08
Qwen2.5-72B-Instruct 9.58 44.36 34.81 17.81 87.13 74.72 21.69
QA MedMCQA Llama-3.1-8B-Instruct 9.70 31.44 18.90 15.01 46.74 29.53 40.35
Qwen3-32B 9.75 49.80 33.27 13.79 65.27 45.36 33.44
Llama-3.1-70B-Instruct 9.95 69.67 49.59 4.52 48.92 32.79 28.90
MMLU Llama-3.1-8B-Instruct 9.99 58.25 37.47 8.47 53.72 33.96 35.72
Qwen3-32B 10.00 82.96 65.22 8.13 78.47 60.40 21.43
Llama-3.1-70B-Instruct 9.96 88.20 72.10 4.18 67.94 52.17 18.24
Models.
We conduct extensive experiments on various open-sourced AI models. For image classification, we utilize three well-established deep image classifiers: ResNet-34 (He et al., 2016), DenseNet-161 (Huang et al., 2018), and ResNext50 (Xie et al., 2017). Additionally, we employ the Vision-Language Model CLIP (Radford et al., 2021), which is based on a Vision Transformer architecture (ViT-B/32) (Dosovitskiy et al., 2021). The above classifiers are provided by TorchVision (Paszke et al., 2019). For text labeling, we employ three LLMs: Llama-3.1-8B-Instruct (Grattafiori et al., 2024), Qwen3-32B, Qwen2.5-72B-Instruct (Qwen et al., 2025). For LLM QA tasks, we employ five LLMs: Qwen3-8B (Yang et al., 2025), Qwen3-14B, Qwen3-32B, Llama-3.1-8B-Instruct, and Llama-3.1-70B-Instruct. The above LLMs are provided by Hugging Face.
Baselines and evaluation metrics.
We evaluate Conformal Labeling with two baseline methods: using AI models to label test instances with uncertainty scores $\mathcal{S}_{n+j}\leq 0.1$, and applying AI predictions to the entire test dataset. We compare Conformal Labeling’s selection procedure against BH (Benjamini & Hochberg, 1995), Storey-BH (Storey, 2002), and Quantile-BH procedures (Benjamini et al., 2006). We evaluate the performance of our method and baseline using the following metrics: (1) FDR, the expected proportion of incorrect labels in the selected set $\mathcal{R}$; (2) Power, the proportion of correctly labeled instances selected; (3) AI-labeled ratio, the number of data labeled by the AI models divided by the combined size of the calibration and test datasets.
Implementation details.
To ensure the reliability of our results, we repeat each experiment 1000 times and report the average result. We randomly select 10% of the data as the calibration dataset. For all the experiments, we use the maximum softmax probability (MSP) as the uncertainty score function. In the LLM QA task, we adopt the standard multiple-choice evaluation pipeline: given a question and candidate answers (e.g., A, B, C, or D), the model estimates the probability of each option by extracting the logits corresponding to the option tokens and applying a softmax transformation, with the predicted label taken as the option with the highest probability. For text labeling, we reformulate each sample into a multiple-choice format (e.g., “positive,” “negative,” or “neutral”), enabling the same probability-extraction procedure as in the LLM QA task. More details of implementation are provided in Appendix F. We provide the code for reproducing our main experiments in this anonymous repository.
4.2 Experiments results
Conformal Labeling achieves tight FDR control with high power.
In table 1, we present the performance of Conformal Labeling against two baselines on three different labeling tasks: image labeling, text labeling, and LLM QA. A salient observation is that across all the labeling tasks and all the model architectures, Conformal Labeling successfully controls the FDR at or below the target FDR level. In comparison, both baseline methods lack FDR control, resulting in substantial labeling errors that compromise label quality when AI models are inaccurate. It is worth noting that the FDR is tightly controlled: for $\alpha=10\%$, most experiments yield FDRs below $9.9\%$, with the largest deviation at $9.56\%$. This tight FDR control directly leads to high selection power. For example, on MMLU with Qwen3-32B at $\alpha=10\%$, Conformal Labeling achieves a power of $82.96\%$, labeling $65.22\%$ of the dataset with the AI model. Overall, empirical results show that Conformal Labeling consistently achieves tight FDR control with high power across different datasets and models.
Figure 2: Performance comparison of Conformal Labeling across models of varying accuracy. We employ Qwen3-8B, Qwen3-14B, and Qwen3-32B (model accuracy increases with parameter count) on MMLU and MMLU-Pro with $\alpha=0.1$. The results show that model with higher accuracy achieves greater power and AI-labeled ratio.
Higher prediction accuracy enables better selection results.
The performance of Conformal Labeling depends heavily on the underlying prediction accuracy, which is influenced by model capacity and dataset difficulty. We evaluate Conformal Labeling with Qwen3-8B, Qwen3-14B, and Qwen3-32B, whose increasing scales provide greater capacity. MMLU-Pro, more challenging than MMLU, results in lower prediction accuracy. Figure 2 presents the evaluation results. In all cases, the realized FDR remains below and close to $\alpha=0.1$. Higher accuracy—achieved with stronger models or easier datasets—boosts power and the AI-labeled ratio. Specifically, for any given model, performance is superior on MMLU compared to MMLU-Pro. Similarly, for each dataset, larger models yield greater power and AI-labeled ratios. Overall, higher prediction accuracy leads to better selection results by achieving higher power and AI-labeled ratio with controlled FDR.
How many calibration samples are needed?
The size of the calibration set plays a crucial role in constructing reliable conformal $p$-values. To study the effect of $|D_{\mathrm{cal}}|$ on FDR and power, in Figure 3, we label {5%, 10%, 20%} of the unlabeled dataset as the calibration dataset. Our results demonstrate that Conformal Labeling is robust to calibration set size: even with a 5% calibration ratio, the FDR remains controlled with low standard deviation. Increasing the calibration ratio reduces the variance of both FDR and power, although the improvement from 10% to 20% is negligible. Based on this trade-off between variance reduction and labeling cost, we use a 10% calibration ratio for all the experiments. In summary, while Conformal Labeling is robust to calibration set sizes, larger calibration sets reduce the variance of FDR and power, thereby enhancing selection stability.
Figure 3: Performance of Conformal Labeling with varying calibration set sizes on the MedMCQA dataset. The top row corresponds to the results from Qwen3-8B and the bottom row to those from Qwen3-32B; each column corresponds to a value of calibration ratio. Shaded regions indicate one standard deviation around the average. The results show that a large calibration set slightly reduces the variance of FDR and power, leading to a more robust selection outcome. Overall, our method is robust to changes in the calibration set size.
Figure 4: Albation study on the selection procedure. The ablation study is conducted on the MMLU dataset with Qwen3-32B. We substitute Conformal Labeling’s selection procedure with alternative selection procedures, including BH, Storey-BH, and Quantile-BH procedures. The figure shows the Conformal Labeling’s selection procedure consistently achieves the tightest FDR control and the highest power among all the selection procedures.
Ablation study on the selection procedure.
To validate the effectiveness of Conformal Labeling’s selection procedure, we conduct an ablation study substituting it with alternative selection procedures, including BH, Storey-BH (Storey, 2002), and Quantile-BH (Benjamini et al., 2006) procedures (see Appendix D for details). The hyperparameters for the Storey-BH and Quantile-BH procedures are chosen using a bootstrap method detailed in the Appendix E. An ideal selection procedure aims to maximize power while keeping the FDR below the desired level.
The results of this ablation study are presented in Figure 4. While all the selection procedures successfully control the FDR below the desired level, Conformal Labeling’s selection procedure (red line) provides the tightest control, with realized FDR consistently closest to the desired level (dashed blue line). In comparison, alternative selection procedures are more conservative, especially at low target FDR levels, which reduces power. Overall, this ablation study highlights the superiority of Conformal Labeling’s selection procedure.
5 Related work
Selection with conformal $p$-values.
Conformal $p$-values are widely studied for their distribution-free, model-agnostic properties in selection tasks. Existing methods fall into two main categories: conformal novelty detection (Bates et al., 2023; Bashari et al., 2023; Wu et al., 2025; Bashari et al., 2025; Lee et al., 2025; Huo et al., 2025) and conformal selection (Jin & Candès, 2023; Gui et al., 2024). Conformal novelty detection aims to identify out-of-distribution instances, while conformal selection aims to select data points that meet a specific quality criterion. For conformal novelty detection, recent advances primarily focus on improving the power of the selection procedure by using various forms of side information (Liang et al., 2024; Marandon et al., 2024; Zhao & Sun, 2025). Jin & Candès (2023) introduced the conformal selection framework for data selection in regression settings, which has since been extended to various settings: multivariate data selection (Bai et al., 2025b), online data selection (Xu & Ramdas, 2024; Liu et al., 2025), and human-in-the-loop adaptive data selection (Gui et al., 2025). Recent works have also applied CS to different tasks such as candidate screening (Lu et al., 2025), drug discovery (Jin & Candès, 2023; Bai et al., 2025a), and foundation model alignment (Gui et al., 2024). However, none of them has studied how to construct conformal $p$-values and design selection procedures in the selective labeling task.
Selective labeling.
Our method builds on the idea of selective labeling, which prioritizes the collection of expert labels for instances where the AI model exhibits uncertainty, and relies on the model’s prediction where it is confident. (Gu et al., 2012; Vrabac et al., 2022). Prior works on selective labeling have explored various heuristic methods, including collaborative annotation frameworks that integrate expert and LLM efforts (Li et al., 2023), and domain-specific applications in text, vision, and medicine (Kim et al., 2024; Vrabac et al., 2022; Duan & Lalor, 2023; Zhang et al., 2025). More recent research aims to provide theoretical foundations, with methods designed to control the overall labeling error (Candès et al., 2025) or to enable valid statistical inference (Zrnic & Candes, 2024; Gligorić et al., 2024). Our method is also closely related to selective prediction, where models are allowed to abstain from making a prediction when uncertain about the output (Geifman & El-Yaniv, 2017; Mozannar & Sontag, 2020; Yang et al., 2023). For LLMs, recent works have also studied how to teach the model not to predict when the model is uncertain (Kamath et al., 2020; Yoshikawa & Okazaki, 2023). These selective prediction methods also lack theoretical guarantees for the AI predictions. Distinct from prior methods, our work is the first to provide a guarantee for the quality of AI-assigned labels, thereby enabling the reliable deployment of selective labeling.
6 Conclusion
In this work, we propose Conformal Labeling, a novel selective labeling method for identifying samples where AI predictions can be provably trusted. This is achieved by controlling the false discovery rate (FDR), which ensures that the expected fraction of incorrect labels in the selected subset is below a user-specified level. The key idea is to reformulate selective labeling as multiple hypothesis testing, which enables distinct theoretical guarantees and methodological advantages compared to prior approaches. In particular, we construct a conformal $p$-value for each test instance by comparing the AI model’s predicted confidence to those of mislabeled calibration instances. Then, we select all the test samples whose $p$-values are smaller than or equal to a data-dependent threshold. We theoretically prove that Conformal Labeling successfully controls the FDR under mild assumptions. Extensive experiments demonstrate that Conformal Labeling achieves tight FDR control and high power across various tasks, including image and text labeling, and LLM QA. We hope the insight from this work will inspire future research to explore reliable selective labeling.
Appendix A Use of Large language model
This paper uses large language models solely to polish specific sentences or paragraphs, without further use of LLMs for other purposes.
Appendix B Technical proofs
B.1 A lemma for proving Theorem 3.1
Lemma B.1.
Under the assumptions of Theorem 3.1, where the calibration samples $(X_{i},Y_{i})_{i=1}^{n}$ and the test samples $(X_{n+j},Y_{n+j})_{j=1}^{m}$ are independently and identically distributed (i.i.d.), and conditional on $n_{0}$ (the number of true null hypotheses in the calibration set) and $m_{0}$ (the number of true null hypotheses in the test set), the expected false discovery proportion (FDP) satisfies
$\mathbb{E}[\text{FDP}\mid n_{0},m_{0}]\leq\frac{1+n}{1+n_{0}}\frac{m_{0}}{m}\alpha,$
where $\alpha\in(0,1)$ is the target false discovery rate (FDR) level, and the selection set $\mathcal{R}$ is determined by Algorithm 1.
Proof.
Suppose the first $n_{0}$ are true null samples in the calibration set and the first $m_{0}$ samples are true null samples in the test set. By the standard result in conformal inference (Vovk et al., 2005), we have
$\mathbb{P}\{\hat{p}_{j}\leq\alpha\mid n_{0},m_{0}\}\leq\alpha,j=1,\ldots,m_{0}$
Let $j^{*}$ denote the number of rejections from Algorithm 1, and recall the false discovery proportion (FDP) as
$\text{FDP}=\frac{\sum_{j=1}^{m_{0}}\mathbf{1}\!\left\{\hat{p}_{n+j}\leq\tfrac{1+n}{1+n_{0}}\cdot\tfrac{\alpha j^{*}}{m}\right\}}{\max(j^{*},1)}.$
Since $\text{FDP}=0$ whenever $j^{*}=0$, we have
$\mathbb{E}[\text{FDP}\mid n_{0},m_{0}]=$ $\mathbb{E}\!\left[\frac{1}{j^{*}}\sum_{j=1}^{m_{0}}\mathbf{1}\!\left\{\hat{p}_{n+j}\leq\tfrac{1+n}{1+n_{0}}\cdot\tfrac{\alpha j^{*}}{m}\right\}\middle|n_{0},m_{0},j^{*}>0\right]\mathbb{P}\left\{j^{*}>0\right\}$ (3)
$\quad\quad+\mathbb{E}\!\left[\sum_{j=1}^{m_{0}}\mathbf{1}\!\left\{\hat{p}_{n+j}\leq\tfrac{1+n}{1+n_{0}}\cdot\tfrac{\alpha j^{*}}{m}\right\}\middle|n_{0},m_{0},j^{*}=0\right]\mathbb{P}\left\{j^{*}=0\right\}$
$\leq$ $\mathbb{E}\!\left[\frac{1}{j^{*}}\sum_{j=1}^{m_{0}}\mathbf{1}\!\left\{\hat{p}_{n+j}\leq\tfrac{1+n}{1+n_{0}}\cdot\tfrac{\alpha j^{*}}{m}\right\}\middle|n_{0},m_{0},j^{*}>0\right]$
$=$ $\mathbb{E}\!\left[\text{FDP}\;\middle|\;n_{0},m_{0},j^{*}>0\right].$
By the super-uniform property of conformal $p$-values,
$\mathbb{E}\!\left[\sum_{j=1}^{m_{0}}\mathbf{1}\!\left\{\hat{p}_{n+j}\leq\tfrac{1+n}{1+n_{0}}\cdot\tfrac{\alpha j^{*}}{m}\right\}\;\middle|\;n_{0},m_{0},j^{*}\right]\leq m_{0}\cdot\frac{1+n}{1+n_{0}}\cdot\frac{\alpha j^{*}}{m}.$
Thus,
$\mathbb{E}\!\left[\text{FDP}\;\middle|\;n_{0},m_{0},j^{*}>0\right]\leq\frac{1}{j^{*}}\cdot m_{0}\cdot\frac{1+n}{1+n_{0}}\cdot\frac{\alpha j^{*}}{m}=\frac{m_{0}}{m}\cdot\frac{1+n}{1+n_{0}}\alpha.$
Combining with (3) yields
$\mathbb{E}[\text{FDP}\mid n_{0},m_{0}]\leq\mathbb{E}\!\left[\text{FDP}\;\middle|\;n_{0},m_{0},j^{*}>0\right]\leq\frac{m_{0}}{m}\cdot\frac{1+n}{1+n_{0}}\alpha,$
which completes the proof. ∎
B.2 Proof of theorem 3.1
Proof of theorem 3.1.
Under the i.i.d. assumption of Theorem 3.1, the calibration samples $(X_{i},Y_{i})_{i=1}^{n}$ and the test samples $(X_{n+j},Y_{n+j})_{j=1}^{m}$ are independently and identically distributed. This implies that each hypothesis, whether from the calibration or test set, has an equal probability $p$ of being a true null, where $p$ represents the expected probability of incorrect prediction under the null hypothesis.
Consequently, the number of true null hypotheses in the calibration set, denoted $n_{0}$, follows a binomial distribution $n_{0}\sim\text{Binomial}(n,p)$, and the number of true null hypotheses in the test set, denoted $m_{0}$, follows $m_{0}\sim\text{Binomial}(m,p)$. The independence across the calibration and test sets arises from the i.i.d. structure, ensuring that $n_{0}$ and $m_{0}$ are independent random variables.
Using the law of total expectation, we express FDR as
$\text{FDR}=$ $\mathbb{E}[\text{FDP}]$ (4)
$=$ $\mathbb{E}[\mathbb{E}[\text{FDP}\mid n_{0},m_{0}]]$
[…]
$=$ $\mathbb{E}\left[\frac{m_{0}}{m}\right]\cdot\mathbb{E}\left[\frac{1+n}{1+n_{0}}\right]\cdot\alpha\quad\text{since }n_{0}\text{ and }m_{0}\text{ are independent}$
$=$ $p\cdot\mathbb{E}\left[\frac{1+n}{1+n_{0}}\right]\cdot\alpha\quad\text{since }m_{0}\sim\text{Binomial}(m,p)$
Now it suffices to show that $p\cdot\mathbb{E}\left[\frac{1+n}{1+n_{0}}\right]=[1-(1-p)^{n+1}]$.
Since $n_{0}$ has probability mass function
$\mathbb{P}(n_{0}=k)=\binom{n}{k}p^{k}(1-p)^{n-k},\quad k=0,1,\dots,n,$
we compute
$p\cdot\mathbb{E}\left[\frac{1+n}{1+k}\right]=$ $p\cdot\sum_{k=0}^{n}\frac{1+n}{1+k}\binom{n}{k}p^{k}(1-p)^{n-k}$ (5)
$=$ $\sum_{k=0}^{n}\frac{(n+1)!}{(k+1)!(n-k)!}\,p^{k+1}(1-p)^{n-k}$
$=$ $\sum_{k=0}^{n}\binom{n+1}{k+1}\,p^{k+1}(1-p)^{n-k}$
$=$ $\sum_{l=1}^{n+1}\binom{n+1}{l}p^{j}(1-p)^{n+1-l}\quad\text{by letting }l=k+1$
$=$ $\sum_{l=0}^{n+1}\binom{n+1}{l}p^{j}(1-p)^{n+1-l}-\binom{n+1}{0}p^{0}(1-p)^{n+1-0}$
$=$ $\sum_{l=0}^{n+1}\mathbb{P}(X=l)-(1-p)^{n+1}\quad\text{where }X\sim\text{Binomial}(n+1,p)$
$=$ $1-(1-p)^{n+1}$
This completes the proof, establishing the desired bound on the FDR.
Appendix C Overview of Different Uncertainty Score Functions
In this section, we provide an overview of three representative uncertainty score functions that are widely used in misclassification detection: Maximum Softmax Probability (MSP) (Hendrycks & Gimpel, 2016), the energy-based score (Liu et al., 2020), and the DOCTOR-$\alpha$ score (Granese et al., 2021). Each of these functions captures predictive uncertainty from a different perspective.
Maximum Softmax Probability (MSP).
The Maximum Softmax Probability (MSP) baseline (Hendrycks & Gimpel, 2016) proposes to use confidence of the AI model as an uncertainty score.
$S_{\mathrm{MSP}}(x)=1-\max_{y\in\mathcal{Y}}p_{y}(x),$
where $p_{y}(x)$ is the softmax probability assigned to class $y$ for input $x$. The key idea is straightforward: if the model assigns a high probability to its most likely class, the prediction is considered confident. Although MSP is simple and effective, it has notable limitations: it only reflects the confidence in the top-1 prediction and ignores the structure of the remaining probability distribution, which may contain useful information about uncertainty.
Energy-based score.
The energy-based score (Liu et al., 2020) is defined as
$S_{\mathrm{Energy}}(x)=\log\sum_{y\in\mathcal{Y}}\exp\!\left(f_{y}(x)\right),$
where $f_{y}(x)$ denotes the logit value for class $y$ and $T>0$ is a temperature parameter. This score is derived from the concept of energy in statistical physics and leverages the log-sum-exp operator over all logits. Unlike MSP, which only considers the maximum probability, the energy score integrates information from the full logit vector, thereby providing a smoother and more informative confidence measure.
DOCTOR-$\alpha$ score.
The DOCTOR-$\alpha$ score (Granese et al., 2021) is defined as
$S_{\alpha}(x)=\sum_{y\in\mathcal{Y}}p_{y}(x)^{2},$
where $p_{y}(x)$ denotes the softmax probability for class $y$. This score is inspired by information-theoretic measures of uncertainty, as it is closely related to the quadratic Rényi entropy. The intuition is that if the predictive distribution is sharp (i.e., one class has probability close to one), then $\sum_{y}p_{y}(x)^{2}$ will be large, indicating high confidence. Conversely, if the distribution is flat (i.e., the model is uncertain and spreads probability mass across many classes), then $S_{\alpha}(x)$ will be small. Compared to MSP, the DOCTOR-$\alpha$ score leverages information from the entire probability distribution rather than only the top prediction, making it a richer measure of uncertainty for misclassification detection.
Appendix D BH Procedure and its Adaptive Variants
Consider testing $m$ null hypotheses $H_{0}^{1},\ldots,H_{0}^{m}$ based on their corresponding $p$-values $\{p_{1},p_{2},\ldots,p_{m}\}$. For a true null hypothesis $H_{0}^{j}$, the corresponding $p$-value $p_{j}$ is a random variable that is super-uniform on $[0,1]$ under the null hypothesis. Formally, for any $u\in[0,1]$,
$\mathbb{P}(p_{j}\leq u\mid H_{0}^{j}\text{ is true})\leq u.$
Define $p_{(j)}$ as the $j$-th smallest $p$-value among a set of $p$-values $\{p_{1},p_{2},\ldots,p_{m}\}$. Given a set of $p$-values $\{p_{1},p_{2},\ldots,p_{m}\}$, the BH algorithm returns $S=\{j\in\{0,\ldots,m\}:p_{j}\leq\frac{\alpha j^{*}}{m}\}$, where $\alpha$ is the target FDR level and
$j^{*}=\text{max}\{j\in\{1,\ldots,m\}:p_{(j)}\leq\frac{\alpha j}{m}\}$
When the null $p$-values $\{p_{j}:j\in\mathcal{H}_{0}\}$ are independent, the BH procedure is proved to control the FDR at level $\pi_{0}\alpha$ in finite samples (Benjamini & Hochberg, 1995), where $\pi_{0}=\frac{|\mathcal{H}_{0}|}{m}$ is the proportion of true nulls. The independence assumption can be further relaxed to the PRDS condition (Benjamini & Yekutieli, 2001).
If $\pi_{0}$ is small, the FDR control will be overly conservative. When $\pi_{0}$ is known, we can apply BH procedure at level $\frac{\alpha}{\pi_{0}}$ to close the gap. In practice, $\pi_{0}$ is typically unknown. Several adaptive BH procedures attempt to address this issue by estimating $\pi_{0}$ and adjusting the target FDR level $\alpha$ accordingly. These procedures are often called the $\pi_{0}$-adaptive versions of the BH algorithm. Two most famous estimators are Storey-BH (Storey, 2002) and Quantile-BH (Benjamini et al., 2006):
$\hat{\pi}_{0}^{Storey}(\lambda)=\frac{1+\sum_{i=1}^{m}\mathbf{1}\{p_{i}\geq\lambda\}}{m(1-\lambda)},\lambda\in(0,1)$
$\hat{\pi}_{0}^{Quant}(k_{0})=\frac{m-k_{0}+1}{m(1-p_{(k_{0})})},k_{0}\in\{1,\dots,m\}$
$\lambda$ and $k_{0}$ are hyperparameters determined by users.
Appendix E Hyperparameter selection for BH adaptive variants
[…]
  1. 1.
    Define a grid $R$ for the hyperparameter, i.e. $R=\{0.1,0.2,\dots,0.9\}$ for Storey-BH.
  2. 2.
    For each $\lambda\in R$, compute:
    $\widehat{\mathrm{pFDR}}_{\lambda}(\gamma)=\frac{\widehat{\pi}_{0}(\lambda)\gamma}{\widehat{\Pr}(p\leq\gamma)\{1-(1-\gamma)^{m}\}}$ (6)
    where $\widehat{\pi}_{0}(\lambda)=\widehat{\pi}_{0}^{Storey}(\lambda)\text{ or }\widehat{\pi}_{0}^{Quant}(\lambda)$ and $\widehat{\Pr}(p\leq\gamma)$ is the empirical estimate of $\Pr(p\leq\gamma)$
  3. 3.
    Generate $B$ bootstrap replicates $\{p_{1}^{*,b},\dots,p_{m}^{*,b}\}_{b=1}^{B}$ and compute $\widehat{\mathrm{pFDR}}_{\lambda}^{*,b}(\gamma)$ for each $b$.
  4. 4.
    Estimate the MSE for each $\lambda$:
    $\widehat{\mathrm{MSE}}(\lambda)=\frac{1}{B}\sum_{b=1}^{B}\left(\widehat{\mathrm{pFDR}}_{\lambda}^{*,b}(\gamma)-\min_{\lambda^{\prime}\in R}\widehat{\mathrm{pFDR}}_{\lambda^{\prime}}(\gamma)\right)^{2}$ (7)
  5. 5.
    Select $\hat{\lambda}=\operatorname*{arg\,min}_{\lambda\in R}\widehat{\mathrm{MSE}}(\lambda)$
Appendix F Implementation details
Experiment details. We run our experiments on NVIDIA GeForce RTX 4090 and NVIDIA L40 GPU, and implement all methods by PyTorch and vLLM.
Dataset details.
For the LLM QA datasets (MedMCQA, MMLU, and MMLU-Pro), we adopt the same prompts as in Luo et al. (2025). For the other datasets, we design our own prompts following the style of LLM QA prompts, as summarized in Table 2. In text labeling and LLM QA tasks, we merge the calibration and test sets to form the unlabeled dataset, except for MedMCQA. For MedMCQA, because test labels are unavailable, we use the calibration set as the unlabeled dataset. For image labeling, we use the validation sets of ImageNet and ImageNet-V2 as the unlabeled datasets.
Table 2: Prompts for different datasets
   Dataset    Prompts    LLM QA    The following are multiple-choice questions. Give ONLY the correct option, no other words or explanation:    [Question] A: [Option 1] B: [Option 2] C: [Option 3] D: [Option 4] Answer: [Mask]        Stance on    global warming    You are given a statement about climate change. Determine the stance that a human would take towards this statement.    Respond with ONLY the letter (A, B, or C) of the correct stance. Do not include any explanation.    [Input Headline] A: agrees B: neutral C: disagrees Answer: [Mask]    Misinformation    You are a fact-checking assistant. Classify the following news headline as either real (A) or misinfo (B).    Respond with ONLY the letter A or B. Do not include any explanation.    Headline: [Input text] A: real B: misinfo Answer: [Mask]
Appendix G Extension to Regression Tasks
While our primary focus is on classification, Conformal Labeling can be naturally extended to regression settings. Consider a loss function L(Y,Y^)L($\hat{Y}$) that quantifies prediction error—for instance, the squared error $L(Y,\hat{Y})=(Y-\hat{Y})^{2}$—alongside a user-specified tolerance level $\epsilon$. For each test sample, we define the null and alternative hypotheses as:
$H_{0}^{j}:L(Y_{n+j},\hat{Y}_{n+j})>\epsilon\quad\text{versus}\quad H_{1}^{j}:L(Y_{n+j},\hat{Y}_{n+j})\leq\epsilon.$
This framework generalizes the classification setting, which corresponds to the special case where $L(Y,\hat{Y})=\mathbf{1}\{Y\neq\hat{Y}\}$ and $\epsilon=0$. To apply Conformal Labeling in regression, we need an uncertainty score that reflects the model’s predictive uncertainty. While uncertainty score functions are straightforward in classification tasks (e.g., $1-\max_{y\in\mathcal{Y}}f_{y}(X)$), regression requires alternative approaches to get an uncertainty score function. For example, when using LLMs, we can leverage their verbalized confidence or prompt them to output prediction intervals, using the interval width as the uncertainty score. The complete procedure for Conformal Labeling in the regression task is outlined in Algorithm 2.
Algorithm 2 Conformal Labeling Regression Tasks
1:Calibration set $\mathcal{D}_{\mathrm{cal}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$; test instances $\{X_{n+j}\}_{j=1}^{m}$; pre-trained predictor $f$; loss function $L$; loss threshold $\epsilon$; target FDR level $\alpha\in(0,1)$; nonconformity scores $\{S_{i}\}_{i=1}^{n+m}$
2:Identify calibration samples that exceed the loss threshold: $\mathcal{D}_{\mathrm{cal}}^{0}=\{(X_{i},Y_{i}):L(Y_{i},\hat{Y}_{i})>\epsilon\}_{i=1}^{n}$, and set $n_{0}=|\mathcal{D}_{\mathrm{cal}}^{0}|$.
3:for $j=1,\dots,m$ do
4:  Compute the conformal $p$-value $\hat{p}_{j}$ according to equation 2.
5:end for
6:Apply the step-up procedure: $j^{*}=\max\left\{j:\hat{p}_{(j)}\leq\frac{\alpha j(n+1)}{m(n_{0}+1)}\right\}$, where $\hat{p}_{(j)}$ is the $j$-th smallest $p$-value.
7:Output: The selected set $\mathcal{S}=\{j:\hat{p}_{j}\leq\hat{p}_{(j^{*})}\}$.
We evaluate Conformal Labeling on two regression problems: sentiment analysis with GPT-4o and protein structure prediction with AlphaFold. We use the data provided by Candès et al. (2025). For sentiment analysis, we prompt GPT-4o to output a prediction interval $a_{i},b_{i}]$ for each target $Y_{i}$, set the predicted value as $\hat{Y}_{i}=(a_{i}+b_{i})/2$, and use the interval length $U_{i}=b_{i}-a_{i}$ as the uncertainty score. For protein structure prediction, we take experimentally derived structures as ground truth and AlphaFold predictions as $\hat{Y}_{i}$. Uncertainty scores are obtained from AlphaFold’s internal confidence measure, the average predicted local distance difference test (pLDDT). For both experiments, we employ the L2 loss function.
In Table 3, we report the performance of Conformal Labeling on regression tasks with $\alpha=0.1$. In all cases, Conformal Labeling consistently controls the realized FDR below the target level. A key observation is that Conformal Labeling’s selection result in regression task is highly sensitive to the choice of the tolerance parameter $\epsilon$. For instance, in sentiment analysis, setting $\epsilon=0.05$ yields an AI-labeled ratio of only $4.55\%$, whereas a slightly larger value $\epsilon=0.06$ increases the ratio to $47.84\%$. In practice, selecting an appropriate $\epsilon$ may require domain expertise or prior knowledge about the task.
Table 3: Performance of Conformal Labeling on regression tasks at $\alpha=0.1$. Results are shown for sentiment analysis (top) and protein folding (bottom) under different values of the tolerance parameter $\epsilon$. In all cases, Conformal Labeling controls the realized FDR below the target level.
Dataset Metric Method CL ($\epsilon=0.05$) CL ($\epsilon=0.06$) CL ($\epsilon=0.07$) Sentiment analysis FDR (%) 8.48% 8.60% 7.89% Power (%) 5.04% 53.14% 94.92% AI-labeled ratio 4.55% 47.84% 85.43% CL ($\epsilon=1$) CL ($\epsilon=4$) CL ($\epsilon=9$) Protein folding FDR (%) 9.67% 9.90% 9.04% Power (%) 27.24% 49.73% 97.90% AI-labeled ratio 10.06% 36.80% 88.00%